OpenStudy (anonymous):

Can I simplify this kp(1-p) / (p^2) into this: k(1-p) / p

OpenStudy (callisto):

yes

OpenStudy (callisto):

you can cross out the p on the top and at the bottom

OpenStudy (anonymous):

are u sure

OpenStudy (anonymous):

yeah

OpenStudy (mertsj):

yes

OpenStudy (callisto):

very sure lol

OpenStudy (anonymous):

thanks lol

OpenStudy (mertsj):

\[\frac{kp(1-p)}{p^2}=\frac{k(1-p)}{p}\]